1. A hydraulic cylinder has bore of 50 mm and moves a mass of 10 kg. It is controlled with a valve with a constant kv = 80 Pa/m. The damping coefficient is 2 Ns/m.
1
Expert's answer
2021-10-26T07:00:16-0400

Find the area:


"A=\pi D^2/4=1.96·10^{-3}\text{ m}^2."

Find the time constant:


"T=\sqrt{\frac{M}{Ak_v}}=7.986\text{ s}."